OpenStudy (anonymous):

What is the simplified form of x^4 - 81/ x + 3?

OpenStudy (anonymous):

Multiply both by the inverse of the denominator (x-3) to get rid of the bottom. That gives you (x-3)(x^4-81) Multiply that out to get x^5-3x^4-81x+243.

OpenStudy (anonymous):

is this the right answer? x^3 + 3x^2 - 9x - 27? @singlesixx

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

I don't know either. But I'm pretty sure it's not the right answer.

OpenStudy (anonymous):

i think its either x^3 - 3x^2 + 9x - 27 x^3 + 3x^2 + 9x + 27 or x^3 + 3x^2 - 9x - 27

OpenStudy (anonymous):

Well, to divide a fraction, you multiply by the inverse of the denominator, which is what I did.

OpenStudy (cwrw238):

you can fist use the difference of 2 squares identity: a^2 - b^2 = ( a - b)(a + b) so x^4 - 81 = (x^2 - 9) (x^2 + 9) and again applying this to first barackets = (x - 3)(x + 3)(x^2 + 9) dividing by ( x + 3): = (x - 3)(x^2 + 9) now expand this

OpenStudy (anonymous):

(x^2 + 9) (x^2 - 9) --- thses are the answer choices. Okay, lets get to solving this problem.

OpenStudy (anonymous):

Okay, I have no idea where that's coming from, so I'm going to go ahead and give up.

OpenStudy (anonymous):

x^3 - 3x^2 - 27 but how is there a 9x in the answer choices?

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

applying the distributive law: (x - 3)(x^2 + 9) = x(x^2 + 9) - 3(x^2 + 9) = x^3 + 9x -3x^2 - 27 = x^3 -3x^2 + 9x - 27

OpenStudy (anonymous):

so basically i had just gave the answer choices which were A - D. I got the answer x^3 - 3x^2 - 27, okay? I don't know how to get the 9x in this answer. Just help me. The closest answer to mine that I got was A, if you look back a little bit above and look at my answer choices. @cwrw238

OpenStudy (anonymous):

okay thanks can you help me with just one more problem, please?

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

Rectangle A has a length of 2x + 6 and a width 0f 3x. Rectangle B has a length of x + 2 and an area of 12 square units greater than Rectangle A's area. What is the simpilifed expression for the width of Rectangle B? @cwrw238

OpenStudy (anonymous):

These are the answer choices, x + 2 x + 1 6x + 6 6(x + 2)(x + 1)

OpenStudy (anonymous):

Ahhhh Hellooooo, I need help. @cwrw238

OpenStudy (cwrw238):

area of rectangle A = length * width = 3x(2x + 6) area of rectangle B = W(x+2) where W is its width so as the difference in areas = 12 sq units:- W(x + 2) - 12 = 3x(2x + 6) now you need to find W in terms of x

OpenStudy (cwrw238):

can you do this ?

OpenStudy (anonymous):

So, how do you simplify it to the next step?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

thats why i need help.

OpenStudy (cwrw238):

ok first add 12 to both sides W(x + 2) = 3x(2x + 6) + 12 now expand RHS w(x + 2) = 6x^2 + 18x + 12 now take out 6 on RHS: W(x+ 2) = 3(x^2 + 3x + 2) next step is to, if possible, factor x^2 + 3x + 2 can u do this?

OpenStudy (cwrw238):

sorry that should be 6(x^2 + 3x + 2) on the 6th line

OpenStudy (cwrw238):

an alternative way would be to divide 6x^2 + 18x + 12 by x + 2 using long division

OpenStudy (anonymous):

i dont know how to factor 6(x^2 + 3x + 2)

OpenStudy (cwrw238):

ok maybe you havent done that in your course have you done long division?

OpenStudy (anonymous):

what is long division? this is simplifying rational expressions.

OpenStudy (anonymous):

okay so Im thinking the answer is D out of my options look above at my answer choices again. @cwrw238

OpenStudy (cwrw238):

No the factors of 6(x^2 + 3x + 2) are 6(x + 1)(x + 2) but but you need to divide this by x + 2

OpenStudy (cwrw238):